# Dicke superradiant enhancement of the heat current in circuit QED

### v1

@@ V Discussion and conclusions
In this work, we have illustrated how and when the collective interactions between $N$ qubits and two thermal baths can enhance the heat transport across the device, compared to a scenario where each qubit is coupled independently with the baths. As we discussed, this device can be implemented within current quantum technologies. In particular, here, we focused on the case of a circuit-QED implementation. We have examined both the parameters relevant to experimental settings and the resilience of the system to a third, parasitic thermal bath that acts locally on each qubit. Our findings indicate that this collective advantage is not only robust but also observable under experimental conditions.
@@
In the future, it will be interesting to investigate similar collective effects when time-dependent driving is present allowing the device to function also as a heat engine Quan et al. (2007); Pietzonka and Seifert (2018) or as a refrigeratorLinden et al. (2010); Mari and Eisert (2012), depending on the specific physical parameters.
@@ VI Acknowledgments
Numerical work has been performed by using the Python toolbox QuTiP2 Johansson et al. (2013) and the PIQS library Shammah et al. (2018). We wish to thank R. Fazio, F. Campaioli, D. De Bernardis, M. Brunelli, D. Ferraro and F. Minganti for useful discussions. G.M.A and M.S. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002955 – CONQUER). P. A. E. gratefully acknowledges funding by the Berlin Mathematics Center MATH+ (AA2-18). F.N gratefully acknowledges funding by the BMBF (Berlin Institute for the Foundations of Learning and Data – BIFOLD), the European Research Commission (ERC CoG 772230) and the Berlin Mathematics Center MATH+ (AA1-6, AA2-8). This work was supported by Research Council of Finland Grant No. 312057 (Centre of Excellence program).
@@ A.1 Quantization of the LC Superconducting Circuit
$\hat{Q}=\sqrt{\frac{\hbar}{2Z_{\rm LC}}}(\hat{a}+\hat{a}^{\dagger})~{},$ (39)
$\hat{\phi}=i\sqrt{\frac{\hbar Z_{\rm LC}}{2}}(\hat{a}^{\dagger}-\hat{a})~{},$ (40)
@@
$\hat{H}_{\text{LC}}=\hbar\omega_{\rm LC}\left({\hat{a}}^{\dagger}{\hat{a}}+\frac{1}{2}\right)~{},$ (41)
where the resonant frequency $\omega_{\rm LC}$ of the LC circuit reads $\omega_{\rm LC}={1}/{\sqrt{LC}}$.
A.2 Quantization of a transmon qubit
@@ A.3 Derivation of the Hamiltonian of the system
where $V_{i}$ is the voltage drop occurring in the resistance $R_{i}$ and $C^{(i)}_{\rm c}$ corresponds to the capacitive coupling between the transmons and the $\rm{RLC}$ circuits. In the following, the capacitive couplings are assumed to be small with respect to the transmon capacities, $C^{(i)}_{\rm c}\ll C_{\rm T}$. The Lagrangian gives rise to canonical momenta $Q_{j}$ (the charges), which are the derivatives of the Lagrangian with respect to $\dot{\phi}_{j}$, $Q_{j}\equiv\partial\mathcal{L}/\partial\dot{\phi_{j}}$. In terms of $\dot{\phi}_{j}$ the charges $Q_{j}$ can be expressed as:
@@
where we neglected terms $\hat{V}^{2}_{i}$ that do not acts on the transmon. By truncating this Hamiltonian on the transmon qubit Hilbert space, as in Eq. (51), we have:
@@ Appendix B Derivation of Lindblad Master Equation
$\hat{H}_{\text{tot}}=\hat{H}+\hat{H}_{\text{RLC}}~{},$ (60)
Here, $\hat{H}_{\text{RLC}}$ represents the Hamiltonian for $RLC$ circuits while $\hat{H}$ is the transmons Hamiltonian defined in Eq. (6). As $\hat{H}_{\text{RLC}}$ contains a resistive part, writing an explicit expression involves the coupling with a collection of infinite $LC$ elements, as discussed for example in Ref. Cattaneo and Paraoanu, 2021. As we will show later, here, we do not need to write down the full $RLC$ Hamiltonian, $\hat{H}_{\text{RLC}}$. Instead, thanks to the fluctuation-dissipation theorem, it will be sufficient to know the classical equation of motion of the $RLC$ circuit. We further define the interaction Hamiltonian $\hat{H}_{\text{int}}$ as:
$\hat{H}_{\text{int}}$ $=\hat{H}_{\text{tot}}-\hat{H}-\hat{H}_{\text{RLC}}~{},$ (61)
@@
We employ the Born-Markov approximation to trace out the field and obtain a master equation governing the evolution of the qubits density matrix $\rho(t)$. The calculations are facilitated by working in the interaction picture. For any generic operator $\hat{O}$, its corresponding form in the interaction picture is $\hat{O}^{\text{I}}(t)=U_{0}^{\dagger}(t)\hat{O}U_{0}(t)$, where $U_{0}(t)=\exp[-i\hat{H}_{\rm bare}t]$ and $\hat{H}_{\rm bare}=\hat{H}_{0}+\hat{H}_{\text{r}}$. Meanwhile, the density matrix in the interaction picture evolves according to $\hat{\rho}^{\text{I}}(t)=U_{0}(t)\hat{\rho}(t)U_{0}^{\dagger}(t)$.
The evolution of the total density matrix in the interaction picture, $\hat{\rho}^{\text{I}}_{\text{tot}}(t)$, is dictated by the following master equation Breuer and Petruccione (2007):
@@
In reaching this equation, we apply the Markov approximation, assuming that the qubits and the resonators remain uncorrelated over the timescales relevant to the evolution of the qubits density matrix. This allows us to approximate $\hat{\rho}^{\text{I}}_{\text{tot}}(t-\tau)\approx\hat{\rho}^{\text{I}}_{\text{tot}}(t)$ under the time integral and to extend the integration limit to $\tau\to\infty$.
@@
The phase factors arise due to the transformation to the interaction picture, where $\omega_{0}$ is the bare frequency of the qubits. By substituting ladder operators in the interaction picture in the master equation, Eq (B) and performing the secular approximation, selecting operators that induce transitions between eigenstates, we obtain
$\partial_{t}\hat{\rho}(t)$ $=$ $-\frac{i}{\hbar}[\hat{H}_{0},\hat{\rho}(t)]-\sum_{i,j}{G_{i}G_{j}}\Big{\{}S_{\hat{V}_{i},\hat{V}_{j}}(-\omega_{0})\big{[}\hat{J}_{+}\hat{J}_{-}\hat{\rho}(t)-\hat{J}_{+}\hat{\rho}(t)\hat{J}_{-}\big{]}+S_{\hat{V}_{i},\hat{V}_{j}}(\omega_{0})\big{[}\hat{J}_{-}\hat{J}_{+}\hat{\rho}(t)-\hat{J}_{-}\hat{\rho}(t)\hat{J}_{+}\big{]}+{\rm h.c.}\Big{\}}~{},$
@@
The voltage dynamical structure factor can be calculated employing the fluctuation-dissipation theorem Cattaneo and Paraoanu (2021):
@@
${\rm Re}[Z_{{\rm tot},i}(\omega)]=\frac{\omega^{2}R_{i}}{\omega^{2}+R_{i}^{2}C^{2}(\omega^{2}-\omega_{\rm LC}^{2})^{2}}~{},$ (72)
This expression can be further simplified by introducing the quality factor $\mathcal{Q}_{i}$ as the ratio of the frequency of the RLC circuit divided by its line width $\kappa_{i}$, $\mathcal{Q}_{i}=\omega_{\rm LC}/\kappa_{i}$. The linewidth of the RLC circuit resonance is given by $\kappa_{i}=R_{i}$. Hence, we have:
${\rm Re}[Z_{{\rm tot},i}(\omega)]=\frac{R_{i}}{1+\mathcal{Q}_{i}^{2}(\frac{\omega}{\omega_{\rm LC}}-\frac{\omega_{\rm LC}}{\omega})^{2}}~{},$ (73)
@@
By plugging the explicit form of the dynamical structure factor in Eq. (B), we can finally arrive at a master equation that can be written in the form of Eq. (9)
@@
$\gamma_{i}=\frac{4\hbar\omega_{0}R_{i}G_{i}^{2}}{1+\mathcal{Q}_{i}^{2}(\frac{\omega_{0}}{\omega_{\rm LC}}-\frac{\omega_{\rm LC}}{\omega_{0}})^{2}}~{}.$ (77)
@@ C.2 Analytical calculation of the heat current
where denote averages over the thermal occupations $P(m_{J}|J)={e^{-m\hbar\omega_{0}/k_{\rm B}T}}/{Z_{J,T}}$ depending on $J$ and on the temperature $T$,
$\braket{\hat{x}}_{J,T}$ $=\text{Tr}\left[\hat{\rho}^{(\rm s)}\hat{x}\right]~{},$ (80)
$=\sum_{m=-J}^{J}\frac{\exp\left(-\frac{m\hbar\omega_{0}}{k_{\rm B}T}\right)}{Z_{J,T}}\braket{J,m}{\hat{x}}{J,m}~{},$ (81)
@@
The our goal reduces in calculating the average value $\langle\hat{J}_{z}\rangle$ in a subspace with a fixed $J$:
@@
Assuming that $J$ is an integer, $Z_{J,T}$ is a geometric series. The sum of this geometric series is given by:
$Z_{J,T}=\frac{\exp\left(\frac{J\hbar\omega_{0}}{k_{\rm B}T}\right)-\exp\left(-\frac{(J+1)\hbar\omega_{0}}{k_{\rm B}T}\right)}{1-\exp\left(-\frac{\hbar\omega_{0}}{k_{\rm B}T}\right)}~{},$ (84)
@@
It is insightful to consider specific limits of Eq. (87). In the regime of low temperatures, characterized by $\hbar\omega_{0}J\gg k_{\rm B}T$, the average $\langle\hat{J}_{z}\rangle_{J,T}$ simplifies to:
@@
This expression is employed in the main text to obtain Eq. (27).
@@
This equation is utilized in the main text to derive Eq. (28).
Additionally, evaluating the ratio ${\langle\hat{J}_{z}\rangle_{J,T}}/({J\langle\hat{J}_{z}\rangle_{1/2,T}})$ in the limit $J\to\infty$ yields:

### v3

@@ V Discussion and conclusions
In this work, we have illustrated how the collective interactions between $N$ qubits and two thermal baths can enhance the heat transport across the device, compared to a scenario where each qubit is coupled independently with the baths.
The device can be implemented within current quantum technologies, and we focus on its circuit-QED implementation. After revealing the physical origin of the collective enhancement of the heat current, we analyze its resilience to a third, parasitic thermal bath that acts locally on each qubit, modeling unavoidable single-qubit noise. We then assess the experimental feasibility of our proposal by choosing system parameters that have been measured in a previous experiment [28], and we propose a minimal experimental device based on two superconducting qubits. Our findings indicate that this collective advantage is not only robust to local noise but also observable under experimental conditions.
@@
Another important limitation of our predictions stems from the use of a Markovian Lindblad master equation, which is well justified only for small transition rates compared to the qubit frequency $\omega_{0}$. This is primarily required by the secular approximation and the weak coupling approximation (commonly referred to as the Born-Markov approximation), as detailed in the derivation of our master equation in Appendix B. Given that we chose relatively small rates, approximately $10\%$ of the bare frequency $\omega_{0}$, the use of the Lindblad master equation is reasonably justified. Additionally, it has been shown that non-secular terms tend to be more relevant during transient dynamics, while their influence is significantly reduced at the steady state [85]. Nonetheless, it would be interesting to explore the influence of non-secular terms and non-Markovian effects that emerge at strong coupling and can play a significant role in an experiment [86].
In the future, it will be interesting to investigate similar collective effects when time-dependent driving is present allowing the device to function also as a heat engine [87, 88] or as a refrigerator [89, 90], depending on the specific physical parameters.
@@ VI Acknowledgments
Numerical work has been performed by using the Python toolbox QuTiP2 [91] and the PIQS library [92]. We wish to thank R. Fazio, F. Campaioli, D. De Bernardis, L. Giacomelli, M. Brunelli, D. Ferraro, and F. Minganti for useful discussions. G.M.A and M.S. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101002955 – CONQUER). P. A. E. gratefully acknowledges funding by the Berlin Mathematics Center MATH+ (AA2-18). F.N gratefully acknowledges funding by the BMBF (Berlin Institute for the Foundations of Learning and Data – BIFOLD), the European Research Commission (ERC CoG 772230) and the Berlin Mathematics Center MATH+ (AA1-6, AA2-8). This work was supported by Research Council of Finland Grant No. 312057 (Centre of Excellence program).
@@ A.1 Quantization of the $LC$ superconducting circuit
$\hat{Q}=\sqrt{\frac{\hbar}{2Z_{LC}}}(\hat{a}+\hat{a}^{\dagger})~{},$ (35)
$\hat{\phi}=i\sqrt{\frac{\hbar Z_{LC}}{2}}(\hat{a}^{\dagger}-\hat{a})~{},$ (36)
@@
$\hat{H}_{\text{LC}}=\hbar\omega_{LC}\left({\hat{a}}^{\dagger}{\hat{a}}+\frac{1}{2}\right)~{},$ (37)
where the resonant frequency $\omega_{LC}$ of the $LC$ circuit reads $\omega_{LC}={1}/{\sqrt{LC}}$.
A.2 Quantization for a transmon qubit
@@ A.3 Derivation of the Hamiltonian of the system
where $V_{i}$ is the voltage drop occurring in the resistance $R_{i}$ and $C^{(i)}_{\rm c}$ corresponds to the capacitive coupling between the transmons and the $RLC$ circuits. In the following, the capacitive couplings are assumed to be small with respect to the transmon capacities, $C^{(i)}_{\rm c}\ll C_{\rm T}$. Under this approximation, in the following, we will trace out the variables $V_{i}$ and obtain a Markovian master equation for the transmons. The Lagrangian gives rise to canonical momenta $Q_{j}$ (the charges), which are the derivatives of the Lagrangian with respect to $\dot{\phi}_{j}$, $Q_{j}\equiv\partial\mathcal{L}/\partial\dot{\phi_{j}}$. In terms of $\dot{\phi}_{j}$ the charges $Q_{j}$ can be expressed as:
@@
where we neglected terms $\hat{V}^{2}_{i}$ that do not act on the transmons. By truncating this Hamiltonian on the transmon qubit Hilbert space, as in Eq. (47), we have:
@@ Appendix B Derivation of Lindblad master equation
$\hat{H}_{\text{tot}}=\hat{H}+\hat{H}_{{RLC}}~{},$ (56)
Here, $\hat{H}_{{RLC}}$ represents the Hamiltonian for $RLC$ circuits while $\hat{H}$ is the transmons Hamiltonian defined in Eq. (6). As $\hat{H}_{{RLC}}$ contains a resistive part, writing an explicit expression involves the coupling with a collection of infinite $LC$ elements, as discussed for example in Ref. 69. As we will show later, here, we do not need to write down the full $RLC$ Hamiltonian, $\hat{H}_{{RLC}}$. Instead, thanks to the fluctuation-dissipation theorem, it will be sufficient to know the classical equation of motion of the $RLC$ circuit. We further define the interaction Hamiltonian $\hat{H}_{\text{int}}$ as:
$\hat{H}_{\text{int}}$ $=\hat{H}_{\text{tot}}-\hat{H}-\hat{H}_{{RLC}}~{},$ (57)
@@
We employ the Born-Markov approximation, which assumes that the qubits are weakly coupled to the modes of the baths, to trace out the field and obtain a master equation governing the evolution of the qubits’ density matrix $\rho(t)$. The calculations are facilitated by working in the interaction picture. For any generic operator $\hat{O}$, its corresponding form in the interaction picture is $\hat{O}^{\text{I}}(t)=U_{0}^{\dagger}(t)\hat{O}U_{0}(t)$, where $U_{0}(t)=\exp[-i\hat{H}_{\rm bare}t]$ and $\hat{H}_{\rm bare}=\hat{H}_{0}+\hat{H}_{\text{r}}$. Meanwhile, the density matrix in the interaction picture evolves according to $\hat{\rho}^{\text{I}}(t)=U_{0}(t)\hat{\rho}(t)U_{0}^{\dagger}(t)$.
The evolution of the total density matrix in the interaction picture, $\hat{\rho}^{\text{I}}_{\text{tot}}(t)$, is dictated by the following master equation [70]:
@@
In reaching this equation, we apply the Markov approximation, assuming that the environment’s correlation time $\tau_{\rm RLC}$ is much shorter than the timescale of the qubit’s dynamics, $\omega^{-1}_{0}$, namely $\omega_{0}\tau_{\rm RLC}\ll 1$. The
This allows us to approximate $\hat{\rho}^{\text{I}}_{\text{tot}}(t-\tau)\approx\hat{\rho}^{\text{I}}_{\text{tot}}(t)$ under the time integral and to extend the integration limit to $\tau\to\infty$.
@@
The phase factors arise due to the transformation to the interaction picture, where $\omega_{0}$ is the bare frequency of the qubits. By substituting ladder operators in the interaction picture into the master equation (Eq. (B)) and performing the secular approximation, we select operators that induce transitions between eigenstates. The secular approximation involves neglecting rapidly oscillating terms in the interaction picture, which is valid when $\omega_{0}$ is large compared to the system-bath couplings. In this regime, we obtain:
$\partial_{t}\hat{\rho}(t)$ $=$ $-\frac{i}{\hbar}[\hat{H}_{0},\hat{\rho}(t)]-\sum_{i,j}{G_{i}G_{j}}\Big{\{}S_{\hat{V}_{i},\hat{V}_{j}}(-\omega_{0})\big{[}\hat{J}_{+}\hat{J}_{-}\hat{\rho}(t)-\hat{J}_{+}\hat{\rho}(t)\hat{J}_{-}\big{]}+S_{\hat{V}_{i},\hat{V}_{j}}(\omega_{0})\big{[}\hat{J}_{-}\hat{J}_{+}\hat{\rho}(t)-\hat{J}_{-}\hat{\rho}(t)\hat{J}_{+}\big{]}+{\rm h.c.}\Big{\}}~{},$ (64)
@@
The voltage dynamical structure factor can be calculated employing the fluctuation-dissipation theorem [69]:
@@
${\rm Re}[Z_{{\rm tot},i}(\omega)]=\frac{\omega^{2}R_{i}}{\omega^{2}+R_{i}^{2}C^{2}(\omega^{2}-\omega_{LC}^{2})^{2}}~{},$ (69)
This expression can be further simplified by introducing the quality factor $\mathcal{Q}_{i}$ as the ratio of the frequency of the $RLC$ circuit divided by its line width $\kappa_{i}$, $\mathcal{Q}_{i}=\omega_{LC}/\kappa_{i}$. The linewidth of the $RLC$ circuit resonance is given by $\kappa_{i}=R_{i}$. Hence, we have:
${\rm Re}[Z_{{\rm tot},i}(\omega)]=\frac{R_{i}}{1+\mathcal{Q}_{i}^{2}(\frac{\omega}{\omega_{LC}}-\frac{\omega_{LC}}{\omega})^{2}}~{},$ (70)
@@
By plugging the explicit form of the dynamical structure factor in Eq. (64), we can finally arrive at a master equation that can be written in the form of Eq. (9)
@@
$\gamma_{i}=\frac{4\hbar\omega_{0}R_{i}G_{i}^{2}}{1+\mathcal{Q}_{i}^{2}(\frac{\omega_{0}}{\omega_{LC}}-\frac{\omega_{LC}}{\omega_{0}})^{2}}~{}.$ (74)
@@ C.2 Analytical calculation of the heat current
where $\braket{\hat{x}}_{J,T}$ denote averages within a subspace with a given $J$, over the thermal occupations $P(m_{J}|J)={e^{-m\hbar\omega_{0}/k_{\rm B}T}}/{Z_{J,T}}$ depending on $J$ and on the temperature $T$,
$\braket{\hat{x}}_{J,T}$ $=\sum_{m=-J}^{J}\frac{\exp\left(-\frac{m\hbar\omega_{0}}{k_{\rm B}T}\right)}{Z_{J,T}}\braket{J,m}{\hat{x}}{J,m}~{},$ (77)
@@
Then our goal reduces to calculating the average value $\langle\hat{J}_{z}\rangle$ in a subspace with a fixed $J$:
@@
This sum can be performed by noticing that $Z_{J,T}$ is a geometric series. The sum of this geometric series is given by:
$Z_{J,T}=\frac{\exp\left(\frac{J\hbar\omega_{0}}{k_{\rm B}T}\right)-\exp\left(-\frac{(J+1)\hbar\omega_{0}}{k_{\rm B}T}\right)}{1-\exp\left(-\frac{\hbar\omega_{0}}{k_{\rm B}T}\right)}~{}.$ (80)
@@
It is insightful to consider specific limits of Eq. (83). In the regime of low-temperatures, characterized by $\hbar\omega_{0}J\gg k_{\rm B}T$, the average $\langle\hat{J}_{z}\rangle_{J,T}$ simplifies to:
@@
This expression proves that the bound in Eq. (28) is tight in the low-temperature regime.
@@
This equation is utilized in the main text to derive Eq. (30). Additionally, evaluating the ratio ${\langle\hat{J}_{z}\rangle_{J,T}}/({J\langle\hat{J}_{z}\rangle_{1/2,T}})$ in the limit $J\to\infty$ yields:
@@
Appendix D Calculation of $\dot{Q}^{\rm det}_{\rm h}$
In this appendix, we detail the calculation of the heat current $\dot{Q}^{\rm det}_{\rm h}$ in the detuned scenario. In this case, each qubit interacts independently with distinct modes of the bath. The total heat current is thus given by the sum of the independent contribution of each qubit, each one determined by Eq. (22) setting $N=1$, with the qubit’s state determined by the master equation in Eq. (33). We assume equal strength in the couplings with both the hot and cold baths. Therefore, in the detuned scenario, the total heat current in Eq. (22) can be expressed as:
$\begin{split}\dot{Q}^{\rm det}_{\rm h}&=\sum_{k=1,2}\hbar\omega_{0}^{(k)}\gamma_{k}\Big{[}-(1+n^{(k)}_{\rm h})\langle\hat{\sigma}^{(k)}_{+}\hat{\sigma}^{(k)}_{-}\rangle+\\ &+n^{(k)}_{\rm h}\langle\hat{\sigma}^{(k)}_{-}\hat{\sigma}^{(k)}_{+}\rangle\Big{]}~{},\end{split}$ (87)
where $\gamma_{k}$ represents the decay rate of the $k$-th qubit, $\hat{\sigma}^{(k)}_{-}$ ($\hat{\sigma}^{(k)}_{+}$) are the creation (destruction) Pauli operators acting on the $k$-th qubit and the thermal occupations $n^{(k)}_{i}$ of the $i$-th bath depend on the bath temperature $T_{i}$ and the frequency of the $k$-th qubit:
$\langle\hat{\sigma}^{(k)}_{+}\hat{\sigma}^{(k)}_{-}\rangle$ $=p^{(k)}_{1}~{},$ (89)
$\langle\hat{\sigma}^{(k)}_{-}\hat{\sigma}^{(k)}_{+}\rangle$ $=p^{(k)}_{0}~{},$
Hence, Eq. (87) can be recasted as:
$\gamma_{2}=\frac{\gamma_{1}}{{1+\mathcal{Q}_{i}^{2}({\omega_{0}}/{\omega_{LC}}-{\omega_{LC}}/{\omega_{0}})^{2}}}~{},$ (92)
with $\mathcal{Q}_{i}$ being the quality factors of the resonators.